OpenStudy (anonymous):

Which function has the flowing behavior? as x approaches ( i'll put the symbol in the comments), y approaches (symbol) as x approaches (symbol), y approaches (symbol) a y=x3-8x2-16 b y=x4-8x2+16 c y=-x4+8x2-16 d y= -x3-8x+16

OpenStudy (anonymous):

Symbol is: \[-\infty -\infty \infty and -\infty\]

OpenStudy (anonymous):

I put the symbols in order according the the question

OpenStudy (johnweldon1993):

So this is asking...as x approaches negative infinity, y approaches negative infinity AND as x approaches infinity, y approaches negative infinity lets test them a) y=x3-8x2-16 if we plug -infinity in here for x.....we also get -infinity for y (-infinity ^ 3 = -infinity) check the other side if we plug in infinity in for x.....we get infinity for y (infinity ^ 3 = infinity) ...so A is not correct. Can you check the others following the same guideline? or would you like more help?

OpenStudy (anonymous):

More help please, I wasn't there for the lesson on this so I have no idea how to do it. :/ sorry if I seem stupid.

OpenStudy (johnweldon1993):

not a problem! and nobody is stupid, you just need to learn it :) so okay...

OpenStudy (johnweldon1993):

We can eliminate B) because we know that -infinity ^ 4th power...would turn that negative...positive...so B cannot satisfy that first behavior . lets check C)...

OpenStudy (anonymous):

Okay. for c it would be the same because it would turn it positive.... right?

OpenStudy (johnweldon1993):

not quite it's important to remember that you do EXPONENTS before multiplication -(-infinity^4) -infinity ^ 4th power does become infinity...however THEN you multiply that by '-' so it DOES become -infinity

OpenStudy (anonymous):

Oh. Okay. I did negative times a negative =positive.

OpenStudy (johnweldon1993):

Right, and don't worry that's a VERY common mistake. It IS true that a negative times a negative = positive...but remember in THIS case....the negative (-) comes AFTER the exponent...now...on with C)...

OpenStudy (anonymous):

Okay.

OpenStudy (johnweldon1993):

C) y=-x4+8x2-16 So we HAVE classified that this satisfies the first behavior....but DOES it satisfy the 2nd? try that...try plugging in infinity for x...and see if you get -infinity

OpenStudy (johnweldon1993):

and remember ...exponents BEFORE multiplication....so its -(infinity^4)

OpenStudy (anonymous):

So yes, it does?

OpenStudy (johnweldon1993):

it does indeed....so since C) satisfies BOTH behaviors...it must be correct....I just want to see if @dumbcow is correcting something I have done incorrectly lol

OpenStudy (anonymous):

Lol. Okay. Thank you for walking me through it. It's a review for my final and it's 80% of my grade.

OpenStudy (dumbcow):

simple way to remember end behavior